# SLogic: Subgraph-Informed Logical Rule Learning for Knowledge Graph Completion

### v1

@@ 5.1 Comparisons with baseline methods
Table 1 shows the effectiveness of SLogic when comparing with the baselines. It shows that SLogic outperforms other methods on WN18RR and YAGO3-10, but performs less effectively on FB15K-237. This is because FB15k-237 contains a larger number of distinct relations, resulting in relatively low rule support and making it harder to distinguish among the rules. The numbers for other systems are taken from Cheng et al. (2022; 2023). Since NCRL employs a different evaluation protocol, where the rank is computed as $m+1$, compared to the strategy used for all other methods. To ensure a fair comparison, we reran NCRL using our evaluation metric and reported these results as NCRL.
@@
None-rule
based
TransE 0.23 2.2 52.4 0.29 18.9 46.5 0.36 25.1 58.0
@@
Rule-based
Learning
AMIE 0.36 39.1 48.5 0.23 14.8 41.9 0.25 20.6 34.3
@@
SLogic 0.49 44.7 55.8 0.28 20.5 42.2 0.50 42.8 63.4
Table 1: SLogic vs. Baselines in KG completion task (Bold/Underlined numbers: best among all methods/best among all rule learning methods; ‘–’: could not be run on our machine)
@@ 5.2 Sensitivity analysis and ablation study
Effect of positive and negative sampling ratio. This analysis evaluates our model’s sensitivity to the two parameters, $k_{\text{pos}}$ and $k_{\text{neg}}$. As shown in Table 2, the results demonstrate that the model is highly robust to different parameter settings. Performance remains stable within a narrow margin across all tested configurations on all the datasets, with the configuration $k_{\text{pos}}=5$ and $k_{\text{neg}}=20$ consistently ranking among the best. The key practical implication of this robustness is that an exhaustive and computationally expensive search for optimal values is unnecessary. This is particularly valuable for larger datasets such as FB15K-237 and YAGO3-10, where memory constraints make more resource-intensive configurations infeasible.
@@
Effect of # of rules per query and subgraph hops. Figure 3 reveals that optimal hyperparameters are highly dependent on the knowledge graph’s structure. As shown in (a), the ideal number of inference rules varies significantly: performance on WN18RR peaks with a moderate 50-70 rules, while YAGO3-10 performs best with the fewest (10), and FB15k-237 shows more volatile behavior. This suggests that for large graphs like YAGO3-10, the top-ranked rules are highly reliable and adding more only introduces noise.
Figure 3(b) shows that a larger local context (more hops) brings more benefit on WN18RR although the performance increase is not obvious when having more than 2 hopes. We cannot access a larger local context using more than $1$-hop for the much denser graphs FB15k-237 and YAGO3-10.
@@ 3 WN18RR 3 FB15k-237 3 YAGO3-10
Figure 3: Sensitivity analysis of MRR across three datasets. The optimal hyperparameters and the feasibility of using deep subgraph contexts vary significantly with KG structure. Higher number of hops could not be run for FB15k-237 and YAGO3-10 on our machine. )
Ablation study and effect of inference hyperparameters. Figure 4 shows the effect of including/excluding the two components of (1) normalizing confidence score and (2) binarization of ground score. Including the normalization score calculation consistently improves the performance (from ✗ to low softmax temperature ($T=0.5$). When this component is included, the performance is not sensitive to the value of $T$. Regarding the binarization component, including it (✗) helps with YAGO3-10, but not the other two datasets.
@@ 5.4 Efficiency analysis
For the rule-based method, we report the time required for rule collection across the different methods as well as the the training time to obtain the performance shown in Table 1. Inference time is excluded from the comparison, since all methods employ comparable inference procedures.
@@
Table 4 shows that SLogic requires more training time than the baseline methods for all the datasets. This overhead arises primarily because generating positive–negative sample pairs substantially increases the number of training instances. Under the default setting of $k_{\text{pos}}=5$ and $k_{\text{neg}}=20$, the resulting number of instances (triplets associated with positive-negative rule pairs) is approximately 100 times larger than the number of triplets in the original graph. For YAGO3-10 dataset, to improve efficiency, only 10% of the graph is used to extract the positive rule base and generate training instances, keeping the training time low. In contrast, identifying negative samples is more time-consuming, as this step requires access to the entire graph.
6 Conclusion
We addressed the problem of knowledge graph completion with a novel rule-based framework. Unlike existing approaches that rely on globally fixed rule confidences, our method leverages the query context, represented as a subgraph of the head entity, to dynamically recalculate rule importance. This context-aware mechanism enables more accurate reasoning by emphasizing rules that are highly relevant to the specific query. Experimental results demonstrate that our approach consistently outperforms other rule-based baselines, as well as embedding-based methods, particularly on knowledge graphs with fewer distinct relations. Beyond improving accuracy, our method also provides interpretable predictions by explicitly grounding query-specific rules.
@@ 6 Conclusion
To facilitate reproducibility, we provide an code repository at: https://anonymous.4open.science/r/slogic-81FE/. The repository contains the datasets used in our experiments, as well as scripts for data preprocessing, model training, and evaluation. It also includes detailed instructions for reproducing all main results reported in the paper.
@@ Appendix B Experimental Setup
B.1 Hardware
@@ B.2 Implementation Details
Appendix C Use of Large Language Models (LLMs)

### v2

@@ 5.1 Comparisons with baseline methods
Table 1 shows the effectiveness of SLogic when comparing with the baselines. It shows that SLogic outperforms other methods on WN18RR and YAGO3-10. On FB15k-237, SLogic achieves competitive results (0.30 MRR), ranking second among rule-based approaches. The performance gap compared to sparse graphs like WN18RR is attributed to the high density of FB15k-237, where valid rules often possess high coverage (reaching many tail entities), introducing ambiguity that dilutes the discriminative power of context-aware scoring. A detailed analysis is in Appendix C.
The numbers for other systems are taken from Cheng et al. (2022; 2023). Since NCRL employs a different evaluation protocol, where the rank is computed as $m+1$, compared to the strategy used for all other methods. To ensure a fair comparison, we reran NCRL using our evaluation metric and reported these results as NCRL.
@@
Non-rule based TransE 0.23 2.2 52.4 0.29 18.9 46.5 0.36 25.1 58.0
@@
Rule-based Learning AMIE 0.36 39.1 48.5 0.23 14.8 41.9 0.25 20.6 34.3
@@
SLogic 0.49 44.7 55.8 0.30 21.9 46 0.50 42.8 63.4
Table 1: SLogic vs. Baselines in KG completion task (Bold/Underlined numbers: best among all methods/best among all rule learning methods; ‘–’: could not be run on our machine). SLogic results are reported as the mean over 5 random seeds with low standard deviation (WN18RR: $0.49\pm 0.0007$, FB15k-237: $0.3\pm 0.0006$, YAGO3-10: $0.50\pm 0.0007$).
@@ 5.2 Sensitivity analysis and ablation study
Effect of positive and negative sampling ratio. This analysis evaluates our model’s sensitivity to the two parameters, $k_{\text{pos}}$ and $k_{\text{neg}}$. As shown in Table 2, the results demonstrate that the model is highly robust to different parameter settings. Performance remains stable within a narrow margin across all tested configurations on all the datasets, with the configuration $k_{\text{pos}}=5$ and $k_{\text{neg}}=20$ consistently ranking among the best. The key practical implication of this robustness is that an exhaustive and computationally expensive search for optimal values is unnecessary. This is particularly valuable for larger datasets such as FB15K-237 and YAGO3-10, where memory constraints make more resource-intensive configurations infeasible. We repeated these sensitivity experiments with 3 random seeds and observed minimal variance ($\sigma\leq 0.002$), confirming that the sensitivity patterns shown in Table 2 are robust to initialization noise.
@@
Effect of # of rules per query and subgraph hops. Figure 3 shows that optimal hyperparameters depend strongly on KG structure. As in (a), the ideal number of inference rules varies widely: WN18RR peaks at 50–70 rules, YAGO3-10 at only 10, and FB15k-237 exhibits more volatile behavior. This suggests that for large graphs like YAGO3-10, the top-ranked rules are highly reliable and adding more only introduces noise. Figure 3(b) shows that a larger local context (more hops) brings more benefit on WN18RR although the performance increase is not obvious when having more than 2 hops. We cannot access a larger local context using more than $1$-hop for the much denser graphs FB15k-237 and YAGO3-10.
@@ 3 WN18RR 3 FB15k-237 3 YAGO3-10
Figure 3: Sensitivity analysis. (Higher number of hops could not be run for FB15k-237 and YAGO3-10 on our machine.)
5.3 Ablation study and effect of inference hyperparameters
Impact of query-dependent scoring. To quantify the benefit of our dynamic scoring mechanism over traditional static approaches, we compared SLogic against a baseline variant, SLogic-Static. This baseline utilizes the identical rule mining and inference pipeline but relies solely on the global Wilson score for ranking, ignoring the query-specific subgraph context.
WN18RRYAGO3-10FB15k-2370.20.20.30.30.40.40.50.50.60.60.370.370.380.380.260.260.490.490.50.50.30.3MRRStatic-Static (Global)SLogic (Dynamic)
(a) Performance comparison: Dynamic vs. Static.
Full ModelNo GNNNo RelEmbNo GRUNo NegSamp0.40.40.450.450.50.50.5030.5030.4550.4550.4350.4350.4280.4280.420.42MRR
(b) Ablation study on YAGO3-10.
Figure 4: Impact of Dynamic Scoring (a) and Component Importance (b). SLogic achieves +17.6% to +25.4% gain over static baselines, and all neural components are essential.
As shown in Figure  4(a), SLogic consistently outperforms the static baseline, confirming the value of context-aware re-ranking. Notably, the performance gains are substantially larger for WN18RR (+24.4%) and YAGO3-10 (+25.4%) compared to FB15k-237 (+17.6%). Our analysis of the rule distributions reveals the cause: on FB15k-237, over 50% of the candidate rules already possess extremely high global confidence (Wilson score >0.9>0.9), leaving little room for improvement via contextual re-scoring. In contrast, WN18RR and YAGO3-10 contain a higher proportion of rules with intermediate confidence (approx. 0.5), scenarios where dynamic, context-specific selection provides the most critical discriminative power.
Ablation study of model components. To validate the necessity of each architectural module, we conducted an ablation study on the YAGO3-10 dataset. We measured the performance impact (MRR drop) when removing four major components one by one.
Figure  4(b) summarizes the results. The significant drop observed when removing Hard Negative Sampling (-16.5%) and the Rule Encoder (-14.9%) indicates that these components form the foundational basis of the model’s reasoning capability. Furthermore, the exclusion of the Subgraph Encoder (GNN) results in a 9.5% performance decrease, confirming that the GNN provides the critical “contextual lift” necessary to push performance beyond static baselines.
Ablation study of inference components. Figure 5 shows the effect of including/excluding the two components of (1) normalizing confidence score and (2) binarization of ground score. Including the normalization score calculation consistently improves the performance (from ✗ to low softmax temperature ($T=0.5$). When this component is included, the performance is not sensitive to the value of $T$. Regarding the binarization component, including it (✗) helps with YAGO3-10, but not the other two datasets.
@@ 5.3 Ablation study and effect of inference hyperparameters
Effect of rule coverage penalty ($\lambda$). On the dense FB15k-237 dataset, introducing the coverage penalty with $\lambda\in[0.5,0.7]$ increased MRR from 0.28 to 0.30 (+7%). This confirms that down-weighting high-coverage rules is critical for dense graphs. (See Appendix C for detailed diagnostic).
@@ 5.5 Efficiency analysis
For the rule-based method, we report the time required for rule collection across the different methods as well as the training time to obtain the performance shown in Table 1. Inference time is excluded from the comparison, since all methods employ comparable inference procedures.
@@
Table 4 shows that SLogic requires more training time than the baseline methods for all the datasets under the default setting. This overhead arises primarily because generating positive–negative sample pairs substantially increases the number of training instances. Under the default setting of $k_{\text{pos}}=5$ and $k_{\text{neg}}=20$, the resulting number of instances is approximately 100 times larger than the number of triplets in the original graph. In contrast, identifying negative samples is more time-consuming, as this step requires access to the entire graph. However, adopting an efficient setting ($k_{pos}=1,k_{neg}=10$) significantly reduces the training time, achieving a $\sim 3.5\times$ speedup on WN18RR and FB15k-237 with negligible MRR drop ($1.3\%$) and a $1.3\times$ speedup on YAGO3-10, offering a practical trade-off for resource-limited scenarios.
6 Conclusion and Discussions
We proposed a novel rule–based learning framework for KGC. Unlike existing approaches that rely on globally fixed rule confidences, our method leverages query contexts (subgraphs) to dynamically recalculate rule importance. This context-aware mechanism resolves uncertainty missed by static scores, enabling more accurate reasoning. SLogic outperforms other rule-based baselines and remains competitive with embedding methods, demonstrating its effectiveness by resolving structural ambiguity and accounting for rule coverage in dense graphs. However, SLogic has limitations. Training is more computationally expensive due to subgraph extraction and hard negative sampling, though sampling reduction offers a workable speed–accuracy trade-off. Finally, SLogic is limited to chain-like Horn clauses; supporting complex non-chain structures and exploring improved graph sampling methods are left for future work.
@@ 6 Conclusion and Discussions
To facilitate reproducibility, we provide an code repository at: https://anonymous.4open.science/r/slogic-81FE/. The repository contains the datasets used in our experiments, as well as scripts for data preprocessing, model training, and evaluation.
@@ Appendix B Experimental Setup
B.1 Default setting of hyperparameters
. For rule base construction, the length of rule body (or the depth of the DFS) $L$ is set to 5 (for WN18RR) and 3 (for both FB15K-237 and YAGO3-10). In subgraph extraction, the number of subgraph hops $k$ is set to be 4 for WN18RR dataset and 1 for both FB15K-237 and YAGO3-10 datasets. The hyperparameter $\alpha$ is 100 for all the datasets. For model training, the margin $\epsilon$ is set to be 1, $k_{pos}$ and $k_{neg}$ are set to be 5 and 20 respectively for all the three datasets. During inference stage, we set the default parameters as $N=50$ for WN18RR and FB15k-237, and $N=10$ for YAGO3-10. For all datasets, we use $T=0.5$ and $\tau=2.0$. Regarding the rule coverage penalty, we set $\lambda=0.65$ for FB15k-237 and $\lambda=0$ for WN18RR and YAGO3-10.
B.2 Hardware
@@ B.3 Implementation Details
Appendix C Impact of Rule Coverage (Diagnostic Analysis)
In Section 5.2, we noted that SLogic performs differently on dense graphs (FB15k-237) compared to sparse ones. To understand this, we analyzed the relationship between the rule scores assigned by SLogic and the rule coverage ($n_{tails}$), defined as the number of distinct tail entities reachable by the rule.
Table 6 shows the average scores assigned to rules grouped by their coverage on FB15k-237. The model assigns significantly higher scores to high-coverage rules.
Rule Coverage ($n_{tails}$) Avg SLogic Score
Low (1-4 tails) -1.12
Medium-Low (4-19 tails) -1.02
Medium (19-73 tails) -0.28
Medium-High (73-339 tails) 0.11
High (339-4123 tails) 0.57
Table 6: Relationship between rule coverage and model scores on FB15k-237.
While high-coverage rules are often valid, they are ambiguous for prediction because they generate hundreds of candidate entities. The high scores assigned to these rules overshadow more specific, discriminative rules. By introducing the penalty term $\lambda\log(n_{tails})$, we explicitly dampen the influence of these high coverage rules. Our experiments confirmed that this penalty yields a 7% performance improvement on FB15k-237, whereas for sparse graphs like WN18RR and YAGO3-10, where rules are naturally specific, $\lambda=0$ remains optimal.
Appendix D Use of Large Language Models (LLMs)